Find the slope of the line 5x-3y=3
Answer by nerdybill(7384)

Put the equation into the "slope-intercept" form:
y = mx + b
where
m is slope
b is the y-intercept
.
5x-3y=3
-3y= -5x + 3
y = (5/3)x - 1
.
Slope is 5/3